Journalist Report – January 9th



Crew 272 Journalist Report 1-9-2023 Sol: 8
Author: Kenny Pritchard, Crew Journalist
Title: Gift Camel

Rejuvenated and well-rested from the previous sol, we began this morning observing a serene, colorful sunrise from the wide window in the Science Dome. Actually, I began the morning by slamming my head into the ceiling while trying to turn off the alarm in my space-conscious room, but that doesn’t sound as good. I threw together some chocolate chip pancakes for the crew after our regular stretching session, and we were all set to go. Adriana covered a brief overview of the EVA as usual, which was planned to be a simple trip to Kissing Camel Ridge to collect geologic samples, observe ancient flow regimes, and test out some more drone scouting.

Kshitij, Mason, Adriana, and I started suiting up for our EVA about a half hour before the planned departure time. Typically, the whole crew heads down to the lower deck to help prepare. The suits are bulky and much easier to don with a helping hand. We gathered all of the equipment we needed, performed radio checks to make sure we’d be able to communicate clearly, adjusted our suit straps as necessary, and stepped into the airlock to prebreathe. Mission Support has presented us with new headsets courtesy of a recent interplanetary cargo shipment, and they worked great! A brief rover drive led us to Kissing Camel Ridge, where we found a surplus of interesting rocks, minerals, fossils, and paleoflow indicators, but a shocking lack of kissing camels. We’re pretty sure we know which formations earned it that name, but it’s still hard to see… Maybe if you squint… Mason did his usual business with the drone and scouted out locations for search & rescue tests he’ll attempt to run in the coming sols. Adriana collected more samples and found some useful info for her research. Kshitij supported Mason’s drone efforts, and I picked up an oversized pocketful of the petrified wood chunks that littered the ground in that area. We finished our objectives and made it back to the Hab with time to spare!

In the GreenHab, Madelyn is in an important phase of her research. She’s pulled some samples of her sprouts to analyze how quickly they’ll wilt after being harvested – part of her investigation is to test claims that the fertilizer she uses will promote longevity and better shelf-life in the plants. She’s also planning for her ultimate harvest on Sol 11 when she will conclude her data collection here at MDRS. Megan was unfortunately forced to scrap a failed hammer print (Pima the 3d printer is a fickle contraption), but luckily, she has already finished printing two others that will be sufficient for the field-testing in a few sols. Lastly, Arly introduced a mid-mission component of her cognitive survey which identifies major contributors to workload on EVA. Because my making breakfast this morning was a schedule adjustment, I also happen to be on dinner tonight, but the crew is generously sparing me from double duty by taking a chunk out of our leftovers instead. Or maybe they just don’t want to eat more of my food? I won’t look a gift horse in the mouth! Or… a gift camel?

Operations Report – January 9th



Crew 272 Operations Report 9-1-2023

SOL: 8

Name of person filing report: Mason Kuhn

Non-nominal systems: Wiring in EVA Suit 7

Notes on non-nominal systems: I have located an issue with the wiring after opening the back of the suit. It could be a damaged wire or a semi-functional soldered connection. I will send a separate email detailing my findings because I have a few attachments I would like Mission Support to look through.


Spirit rover used: No

Hours: 206.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 111.5

Beginning charge: 100%

Ending charge: 67%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 219.0

Beginning charge: 100%

Ending charge: 64%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.5

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: Opportunity and Curiosity were used on today’s EVA. Nominal performance from both.

Summary of Hab operations: Per Mission Support’s request during 272’s mission training, I have replaced the zip-tie connections on the 2 piece suits (from the backpack to the helmet collar) with the provided steel cable ties. I tried on the suit to test its comfortability, and it was similar to slightly more comfortable (as well as more durable). I also opened the backs of suits 7 (non-nominal) and 9 (nominal) to compare the two electrical wiring systems. An issue was located in suit 7’s wiring.

WATER USE: 22.9 gallons

Water (static tank): 226.5 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: Yes

Summary of internet: Nominal

Summary of suits and radios: Suits 1, 5, 9, and 10 were used during the EVA to Kissing Camel Ridge. All suits performed nominally with airflow and minimal fogging. All suits have been cleaned and are currently charging. There were no issues with radios or headsets on the EVA, and the new headsets are an improvement! All radios that were used have been charged and removed from the charging dock.

Summary of GreenHab operations: Monitored and watered GreenHab crops and experimental crops. The lighting system was found to be unplugged, and it was plugged back in last night by mission support. The lights did turn on last night and are fully functional.

WATER USE: 7 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 6g beans, 30g dill

Summary of Science Dome operations: No crew activity, there was Mission Support activity.

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: N/A

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support:

  • Please see my next email for information about suit 7

Best Regards,

Mason Kuhn

Crew 272 Engineer

EVA Report – January 9th



Crew 272 EVA Report 09-01-2023

EVA # 6

Author: Arly Black

Purpose of EVA: 1) First geology EVA to Kissing Camel Ridge. Specific focus on evidence of KCR as an inverted paleochannel as proposed by Clarke and Stoker (2011). 2) Geologist vs non-Geologist field performance evaluation. 3) Drone scouting of surrounding area for feasible search and rescue simulation location of future EVA.

Start time: 11:00 AM

End time: 2:00 PM

Narrative: Upon placing the fancy, new, plush headsets on their heads, the EVA crew knew their days of comms issues were over. It was love at first “copy”. With the previous headsets banished to the town of non-nominal toys, EVA prep went off without a hitch, and all explorers were suited up, thrust into the airlock, and exposed to awful puns with time to spare. The crew exited the airlock and danced their way to Opportunity and Curiosity (don’t think we didn’t see those moves Kenny and Kshitij). They drove south along Cow Dung Road, parked where the road intersects with Kissing Camel Ridge, and continued on foot 100 m south of the west ridge. Kshitij and Mason proceeded to use the drone to scout out potential locations in which to perform their search and rescue research operations. They found that the southern part of the ridge is simply perfect for their planned activity due to clear lines of sight and several good hiding places for a lost astronaut. The drone range reached 2800 ft, more than enough to cover the search grid of interest. Mason then took videos of crew members enjoying their time out in the field.

Simultaneously, Adriana and Kenny went further up the ridge, but within light-of-sight, searching for more paleoflow indicators (Adriana is worried about overusing those words, but I don’t think I’ll ever get sick of them). She took dip measurements, documented around 12 different sites of paleoflow indicators, and wrote some lithologic descriptions. She collected samples with sedimentary structures in them and excitingly, she believes she has found coal seams (geologists rejoice!). Along with minor frolicking and rock-hugging, Kenny documented crew activities and participated in further geologist vs non-geologist performance tests. Besides formations allegedly shaped like large, amorous, humped mammals, the ridge is also home to a staggering amount of petrified wood, which the crew gathered in droves! After taking a final group shot with the drone, the team packed up their supplies and their hard-earned booty and made it back to the Hab a half hour early as they had achieved all their objectives. Ultimately, this was a very successful EVA and the remaining crew back at the Hab can’t wait to get out there to collect some colourful terrified tree bark of their own!

Destination: Kissing Camel Ridge

Coordinates (use UTM WGS 84): N518400, E4249500

Participants: Kshitij Mall (Crew Commander), Mason Kuhn (Crew Engineer), Kenny Pritchard (Crew Journalist), Adriana Brown (Crew Geologist)

Road(s) and routes per MDRS Map: Drove south on Cow Dung Road until Kissing Camel Ridge. Walked on foot along ridge to the east and west relative to Cow Dung Road. Returned to rovers on foot and drove back to the Hab via Cow Dung Road.

Mode of travel: Driving and walking

GreenHab Report – January 9th



Crew 272 GreenHab Report 09-01-2023

GreenHab Officer: Madelyn Whitaker

Environmental control: heater, fan

Average temperatures: 67 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 6 gallons

Daily water usage for research and/or other purposes: 1 gallon

Water in Blue Tank : 243

Time(s) of watering for crops: 1000, 1600

Changes to crops: Harvested beans and dill.

Narrative: Monitored and watered GreenHab crops, monitored and watered experimental crops. Harvested beans for a snack, dill for dill bread.

Harvest: (include which crop and mass in grams): 6g beans, 30g dill

Support/supplies needed: Lighting system was found to be unplugged. Plugged in last night by Dr. Rupert. Lights turned on last night, is now fully functional.

Sol Summary Report – January 9th



Crew 272 Sol Summary Report 09-01-2023

Sol: 8

Summary Title: Rendezvous with the Kissing Camel Ridge

Author’s name: Kshitij Mall

Mission Status: All nominal

Sol Activity Summary: After getting well rested and watching one-third of the Bollywood movie Lakshya (Objective), the crew woke up early to see a beautiful Martian sunrise. HSO Megan led the morning exercise session today for warming up the crew. The Mission Support approved of the EVA for the sol, so the crew was excited to explore another side of Mars: Kissing Camel Ridge. After a delicious Choco-chip pancake breakfast prepared by Crew Journalist Kenneth assisted by me, the crew was briefed on the EVA for the sol. The mission was simple yet important for the Crew Geologist Adriana to study paleo currents in the Martian world. Adriana was supported in this EVA by me, Crew Engineer Mason, and Kenneth. Mason and I also planned to use this EVA to scout sites at the Kissing Camel Ridge area where a Search and Rescue (SAR) operations can be safely and easily conducted. After being helped onsite by Hab crew comprising the Greenhab Officer Madelyn, XO Arly, and Megan. Oppy and Curie were our rovers for the beautiful expedition to the Kissing Camel Ridge.

Upon reaching the Kissing Camel Ridge, Adriana noted some sites for her research to be done in her upcoming EVAs. Mason and I conducted two brief scouting missions using Garud the drone, on the East side of the Kissing Camel Ridge, to identify areas of interest for Adriana and to also look for safe and simple locations to do a simple SAR activity in the upcoming EVA. Along with Kenneth, Mason and I were also asked some simple questions regarding geological activity around the west side of the Kissing Camel Ridge. Our answers as a non-geologist did satisfy the expert geologist on our crew. Adriana and the remaining crew also collected a lot of rock samples to bring back to the Hab. We return 30 minutes prior to our requested EVA return time. A big shoutout to the Mission Support for getting the crew new radio headsets shipped to us by interplanetary cargo ship.

After returning from the EVA, everyone had lunch and filled in the personal survey and completed the perception tests requested by Arly. The EVA crew also completed an additional survey requested by Arly pertaining to human performance during an EVA. Madelyn also tested the shelf-life and welting of the samples. Madelyn was kind to the team by getting some green beans and preparing Dill garlic bread for the crew. PIMA, our 3D printer, had a hiccup printing a big rock hammer for our future expedition. Megan is relooking in the 3D printing setup and got a rock hammer already ready for field testing. Adriana and I also shared with the crew about all the upcoming planned EVAs and took everyone’s opinions to finalize the remaining mission plan. Tomorrow, we plan to conduct our first SAR operation and run a comparison between the effectivity of humans vs. drones for such SAR operations. The dinner for tonight is mostly the leftovers and a pot of ramen prepared by Kenneth and assisted by Megan. Mason and I are also troubleshooting one-piece suit #7 tonight. We plan to get a good night rest and have a successful SAR activity tomorrow.

Look Ahead Plan: Conduct first simple Search and Rescue activity comprising human vs. drones south of Kissing Camel Ridge

Anomalies in work: None

Weather: Overcast, Winds Calm

Crew Physical Status: Good

EVA: One

Reports to be filed: Sol Summary, Operations Report, Journalist Report, EVA Report, EVA Request, Greenhab Report.

Support Requested: None

Sol Summary – December 28th

Sol: 10

Summary Title: Down day

Author’s name: Marc Levesque, Commander

Mission Status: Nominal

Sol Activity Summary: Because of uncertain Martian weather, the crew took a day off from its usual project activities to recoup and prepare for the end of the mission. Games, exercise, harvesting, and naps were among the various activities throughout the sol as we awaited a favorable weather report for tomorrow.

Look Ahead Plan: Two EVAs are planned. The morning will see a reconnaissance of the North Ridge to scout for a possible route up the ridge and site for relocating the radio repeater as part of Levesque’s project. A second afternoon EVA will explore the area around Summerville for Guariniello’s geology project.

Anomalies in work: None

Weather: Cloudy, damp, and cool

Crew Physical Status: Good

EVA: None

Reports to be filed: Sol Summary, Operations, Journalist, GreenHab, Astronomy, and EVA requests.

Support Requested: None

GreenHab Report – December 28th

Crew 271 GreenHab Report 28-12-2022
GreenHab Officer: Alicyn Grete
Environmental control: heater
Average temperatures: 68 degrees F, about 20 degrees C
Hours of supplemental light: 4 hours
Daily water usage for crops: 13 gallons
Daily water usage for research and/or other purposes: 0 gallons
Water in Blue Tank 105 gallons
Time(s) of watering for crops: 1200, 1630
Changes to crops: Harvested 35g beans, 125g microgreens, 58g radishes, 65g dill, 110g carrots, 5g chives, 5g basil
Narrative: This morning I watered all plants and harvested many things. I harvested all the microgreens in the planter by the water tank, then replanted them with 2 packs of umami microgreens (picture attached). Next, I harvested radishes in the planter with the beans and replanted that area with more beans. Then, I harvested dill and carrots to thin those areas out, and harvested the tops of the chives. Later this evening, I harvested basil for our pizza and watered the new plots again.
Harvest: 35g beans, 125g microgreens, 58g radishes, 65g dill, 110g carrots, 5g chives, 5g basil
Support/supplies needed: If you have the time, please check the new things I planted to make sure they look good!

Operations Report – December 28th

SOL: 10

Name of person filing report: Sergii Iakymov

Non-nominal systems: Kitchen sink cabinet hinge, RAM roof window
Notes on non-nominal systems: 1) previously suggested repair on the sink cabinet hinge did not work. It is possible that we must use fasteners to hold it in place. 2) RAM roof window leaked today when it rained/snowed.

Spirit rover used: No
Hours: 206.0
Beginning charge: Not measured
Ending charge: N/A
Currently charging: Yes

Opportunity rover used: No
Hours: 110.4
Beginning charge: Not measured
Ending charge: N/A
Currently charging: Yes

Curiosity rover used: No
Hours: 217.8
Beginning charge: Not measured
Ending charge: N/A
Currently charging: Yes

Perseverance rover used: No
Hours: 252.5
Beginning charge: Not measured
Ending charge: N/A
Currently charging: Yes

General notes on rovers: None

Summary of Hab operations:
WATER USE 19.61 gallons
Water (static tank): 282.07 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes
Summary of internet: Nominal
Summary of suits and radios: Hip belt buckle of suit #11 transferred to suit #5. Broken buckle marked and put away into the cabinet with EVA suits spare parts.

Summary of GreenHab operations:
WATER USE: 13 gal
Heater: On
Supplemental light: plugged
Harvest: Harvested 35g beans, 125g microgreens, 58g radishes, 65g dill, 110g carrots, 5g chives, 5g basil

Summary of ScienceDome operations: used for the Eifert project.
Dual split: On
Summary of RAM operations: N/A.
Summary of any observatory issues: N/A.
Summary of health and safety issues: None
Questions, concerns and requests to Mission Support: N/A.

Copyright © The Mars Society. All rights reserved. | Main Site